### v1

# Stronger together? The homophily trap in networks

Abstract.
While homophily—the tendency to link with similar others—may nurture a sense of belonging and shared values, it can also hinder diversity and widen inequalities. Here, we unravel this trade-off analytically, revealing homophily traps for minority groups: scenarios where increased homophilic interaction among minorities negatively affects their structural opportunities within a network. We demonstrate that homophily traps arise when minority size falls below 25% of a network, at which point minority homophily comes at the expense of lower structural visibility for the minority group. Our work reveals that social groups require a critical size to benefit from homophily without incurring structural costs, providing insights into core processes underlying the emergence of group inequality in networks.
Introduction
Homophily is ubiquitous—people tend to associate with similar others in different settings of social life, from education to relationships to employment (?, ?, ?). This preference for in-group ties is a trade-off: it strengthens social groups through segregation. While homophily nurtures a sense of belonging, it can also limit the access to out-group opportunities and exacerbate inequalities, particularly in networks with minority groups (?, ?, ?). Although this trade-off is a core building block of social networks, it remains poorly understood and analytically unexplored.
By connecting with similar others, people skew their networks towards shared traits, reinforcing social identities and fostering a strong sense of community that provides advantages at the individual level. For example, immigrants settling in an unfamiliar country often leverage intra-ethnic contacts to find economic opportunities (?, ?, ?) and access emotional support networks that are less available across ethnic boundaries (?). This so-called in-group favoritism can further create ideal conditions for collective action. In environmental activism, for instance, activists are more likely to engage in actions when they strongly identify and interact with their groups (?). Such a preference for the familiar, however, inherently hinders diversity by segregating social groups.
This segregation can deprive individuals of accessing opportunities and participating in decision-making processes (?, ?), becoming especially problematic in the workplace, where homophily affects recruitment directly (?, ?). For example, job referrers are more likely to recommend candidates of the same gender, race, age, and education (?, ?, ?), which leads to in-group favoritism that explains occupational segregation (?). This group preference can also contribute to a gender pay gap in healthcare, as physicians tend to refer patients to specialists of the same gender (?). Beyond the workplace, group segregation can shape how individuals perceive others depending on group affiliation, producing perception and implicit biases in networks (?, ?). Overall, while homophily offers benefits, it entails inherent costs.
Such a trade-off is particularly critical in networks with minorities. When minority members favor in-group ties, they inherit not only the opportunities but also the limitations of their group. For instance, at social gatherings, homophily within small social groups limits individuals’ contact pool, resulting in fewer connections on average for these numerical minorities (?). These inherited limitations may also stem from the social capital of a group. For example, immigrants relying on intra-ethnic contacts to find jobs might end up in low-wage positions, potentially leading to an ethnic mobility trap that hinders upward social and economic mobility (?, ?). Nevertheless, despite the negative effects on minority groups, this trade-off remains poorly understood, lacking an analytical approach to identify its underlying mechanisms and understand how intrinsic structural limits in networks sustain it.
In this work, we explore homophily in networks analytically to disentangle its inherent trade-off, with a focus on homophily among minorities. We investigate when homophilic ties are detrimental to minority groups, introducing the concept of the homophily trap—scenarios where increased homophilic interaction among minorities negatively affects their structural opportunities within a network. To study these scenarios, we use a generative network model to construct networks of different group mixing and minority sizes. We show that homophily traps arise when the minority group size falls below 25% of a network. Below this threshold, higher homophily within the minority group leads to fewer structural opportunities for the group: in-group ties come at the expense of lower structural visibility. This trade-off holds regardless of the homophily level in the majority, making it difficult for numerical minorities to both maintain a high number of connections and belong to homophilic social groups. By disentangling the trade-off of homophily analytically and systematically, we build a foundation for understanding how homophily shapes structural opportunities in networks.
Results
To explore homophily and its implications, we focus on relational properties that affect how social groups interact in networks. Specifically, we are interested in relational properties $h_{ij}$ that encode the mixing tendency between groups, which reflects processes such as choice homophily (i.e., ties due to individual preferences), induced homophily (i.e., ties from homogeneous opportunities such as schools or workplaces), and higher-order effects (?, ?). While individual-level properties, such as node degree or fitness, can affect tie formation, they have minimal impact on group-level interaction when assuming independence. Under this assumption, $h_{ij}$ determines the probability of a node from group $i$ to connect with a node from group $j$. For two groups, 00 and $1$, we define $h_{00}$ and $h_{11}$ as the intra-group connection tendencies, with the inter-group tendencies given by $h_{01}=1-h_{00}$ and $h_{10}=1-h_{11}$.
We quantify access to structural opportunities using the average degree of the nodes within each group, given that the number of connections relates to social capital (?, ?), and having fewer connections can result in structural disadvantages (?, ?). We determine the average group degrees by first finding the inter- and intra-group edge counts, which depend on relational properties and group sizes. We find that, on average, the intra-group edge counts are given by $E_{00}=h_{00}n_{0}(n_{0}-1)/2$ and $E_{11}=h_{11}n_{1}(n_{1}-1)/2$, where $n_{i}$ represents group size (see Methods for inter-group edges). With these edge counts, we determine the average group degrees $\langle k_{0}\rangle$ and $\langle k_{1}\rangle$, expressed as $\langle k_{i}\rangle=(2E_{ii}+E_{ij}+E_{ji})/{n_{i}}$.
We use this expression to examine how group mixing affects average group degrees, focusing on scenarios where group 00 is a numerical minority (i.e., $n_{0}<n_{1}$). By varying group sizes and relational properties, we can explore the impact of increasing minority homophily on average degrees. For example, we find that when the minority makes up 20% of the network, higher homophily decreases its average degree, whereas at 30%, higher homophily increases it (see Fig. 1a-b), indicating that the interplay of group mixing and size exhibits a transition state where homophily has double-edged sword effects. This interplay can limit minority structural opportunities and exacerbate a degree gap among groups.
{adjustwidth}
-.25in0in Refer to caption
Figure 1: The homophily trap: when favoring intra-group ties incurs structural costs. (a, b) The impact of higher minority homophily on degree gap depends on group size. (c) When the minority size is 30% of the network, higher minority homophily benefits the group; at 20%, however, it becomes detrimental to the group. (d) We examine this impact across different minority sizes and majority homophily values, finding that with larger minority groups, the impact of homophily shifts from increasing to decreasing the gap. (e) This shift occurs at a critical size, beyond which homophilic interaction helps the connectivity of the minority group. In large networks, scenarios of homophily trap occur when a group is smaller than 25%, where favoring intra-group ties incurs reduced connectivity. In all plots, dots represent simulation results for networks with 1000 nodes, while curves are based on analytical formulations.
Group-size dependence on gaps in structural opportunities.
To investigate the role of homophily in creating structural gaps, we examine analytically how the impact of minority homophily varies with group size. First, we define structural gap $\Delta k=\langle k_{0}\rangle-\langle k_{1}\rangle$ to represent the difference in structural opportunities between the minority and majority groups; then, we examine the impact of higher minority homophily on this gap, revealing that its effect depends on group size (Fig. 1c). Specifically, increasing minority homophily can either widen or narrow the gap, depending on the relative minority size $f_{0}=n_{0}/N$. For small minority fractions (e.g., 20% of the network), higher minority homophily widens the degree gap, limiting the minority group structurally.
We investigate this group-size dependence systematically to characterize the interplay between homophily and structural gaps in networks. We analyze the impact of a fixed increase in minority homophily on the degree gap across various minority fractions and majority homophily values. Our results reveal that as the minority group grows larger, the impact of a higher minority homophily shifts from widening to narrowing the gap (Fig. 1d). Regardless of majority homophily values, a higher minority homophily increases the gap when the minority is small, whereas with a larger minority (e.g., 30% of the network), the gap decreases with higher homophily. This shift represents a qualitative change in the impact of homophily on structural opportunities, depending on the size of the minority group.
Homophily traps in networks with numerical minorities.
To understand how minority homophily shifts from increasing to decreasing structural gaps, we examine the gap analytically to identify the critical minority size when the shift occurs. Precisely, we analyze how the gap changes with variations in minority homophily through examining the derivative of the gap $\Delta k$ with respect to $h_{00}$, leading to the following expression:
$\dfrac{\partial\Delta k}{\partial h_{00}}=2Nf_{0}-\dfrac{N}{2}-1,$ (1)
which describes how the degree gap between minority and majority groups responds to changes in minority homophily. This derivative reveals that the effect of minority homophily on the gap depends on the size of the minority group and the overall network size but not directly on the homophily within the minority or majority groups; that is, the derivative is independent of $h_{00}$ and $h_{11}$.
We are particularly interested in when this derivative equals zero, as it reveals the critical minority size $f_{0}^{*}$ when the impact of minority homophily on the gap shifts from positive values to negative values. By setting Eq. (1) to zero, we find that the critical minority size is given by:
$f_{0}^{*}=\dfrac{2+N}{4N}.$
This critical size depends on the network size $N$, but in larger networks, it becomes independent of $N$ and approaches a fixed value:
$\lim_{N\to\infty}\dfrac{2+N}{4N}=\dfrac{1}{4}=0.25,$
meaning that in sufficiently large networks, the critical minority size is 25% (Fig. 1e). We validate this result through simulations, confirming the existence of $f_{0}^{*}$ in different mixing scenarios and network sizes.
When the minority group is smaller than this critical size, higher minority homophily results in wider structural gap in opportunities, leading to a case of homophily trap, which occurs due to inherent structural limits in the network. A small minority group lacks the size to create enough intra-group edges to compensate for the inter-group edges that are lost with high group homophily. Although higher homophily fosters a more close-knit minority group, these structural limits ultimately reduce broader network access for the group.
Discussion
While homophily fosters in-group cohesion, it implies reduction in out-group interaction, restricting individuals’ access to opportunities beyond their social groups. We disentangle this trade-off, focusing on how homophily among minorities affects their structural opportunities in networks. We show that when a minority group is smaller than 25% of a network, its intrinsic structural limits prevent it from both maintaining a homophilic group and attaining a high number of connections—an instance of the homophily trap. This is independent of the degree of homophily in the majority group.
The homophily trap depicts when favoring intra-group ties incurs structural costs inherited from intrinsic limitations in the group. Our work focuses on limitations regarding group size and access to opportunities, demonstrating that, in small groups, homophily comes with the cost of reduced structural opportunities, as measured via average group degree. This scenario of the homophily trap stems from the natural limits of small groups lacking sufficient members to form homophilic ties. Such a constraint translates into the existence of a necessary critical mass for the minority group to prevent these structural costs.
Similar critical mass thresholds have been qualitatively predicted and empirically observed in different social contexts. In leadership representation, qualitative studies suggest that 30% is the critical proportional number of women required to challenge existing social conventions (?, ?). This threshold aligns with experimental research on social coordination, which has shown that persistent minorities can overturn established social conventions when they reach a critical size—approximately 25% of the population (?). This value for critical mass has been consistently observed in simulation and empirical studies across disciplines, marking a tipping point where persistent minorities can instigate social change (?).
Overall, our work advances the understanding of homophily through an analytical lens on its trade-off. We introduce the concept of the homophily trap to describe scenarios where this trade-off skews toward structural disadvantages for minority groups. We show that homophily traps can arise from structural limits alone, even in networks without mechanisms such as preferential attachment or triadic closure, providing a baseline for future exploration of how network topologies and additional mechanisms affect the homophily trade-off. By disentangling this trade-off, we build a foundation for understanding how homophily shapes opportunities and underpins group inequality in networks.
Methods
To explore homophily and its implications, we focus on the relational properties $h_{00}$ and $h_{11}$, which represent the tendencies for intra-group connections within each group 00 and $1$. We define the inter-group relational properties as $h_{01}=1-h_{00}$ and $h_{10}=1-h_{11}$, ensuring that the probabilities sum to one.
To examine the degree gap between the two groups, we first calculate the average degree for each group. We express the number of nodes in each group as $n_{0}=f_{0}N$ and $n_{1}=(1-f_{0})N$, where $f_{0}$ represents the relative size of the minority group. By using these definitions, we find that, on average, the inter- and intra-group edges counts can be expressed as:
$E_{00}=\dfrac{n_{0}(n_{0}-1)h_{00}}{2},\ E_{01}=\dfrac{n_{0}n_{1}h_{01}}{2},\ E_{10}=\dfrac{n_{1}n_{0}h_{10}}{2},\;\text{and}\;E_{11}=\dfrac{n_{1}(n_{1}-1)h_{11}}{2}.$
With these edge counts, we determine the average degree for each group:
$\langle k_{0}\rangle=\dfrac{2E_{00}+E_{01}+E_{10}}{n_{0}}\quad\text{and}\quad\langle k_{1}\rangle=\dfrac{2E_{11}+E_{10}+E_{01}}{n_{1}}.$
Finally, we define the structural gap between groups as the difference between average degrees:
$\Delta k=\langle k_{0}\rangle-\langle k_{1}\rangle.$
Next, we explore how this structural gap changes with variations in minority homophily, $h_{00}$. To accomplish this, we take the derivative of $\Delta k$ with respect to $h_{00}$:
$\dfrac{\partial\Delta k}{\partial h_{00}}=\dfrac{\partial\langle k_{0}\rangle}{\partial h_{00}}-\dfrac{\partial\langle k_{1}\rangle}{\partial h_{00}}.$ (2)
First, we differentiate each term separately:
$\dfrac{\partial\langle k_{0}\rangle}{\partial h_{00}}=(n_{0}-1)-\frac{n_{1}}{2}\quad\text{and}\quad\dfrac{\partial\langle k_{1}\rangle}{\partial h_{00}}=-\frac{n_{0}}{2}.$
Then, we substitute these expressions into Eq. (2) and replace $n_{1}$ with $N-n_{0}$, yielding:
$\dfrac{\partial\Delta k}{\partial h_{00}}=n_{0}-1-\frac{N-n_{0}}{2}+\frac{n_{0}}{2}=\frac{4n_{0}-N-2}{2}.$
Given that $n_{0}=Nf_{0}$, we can rewrite this expression as:
$\dfrac{\partial\Delta k}{\partial h_{00}}=2Nf_{0}-\dfrac{N}{2}-1.$
To find the minority critical size, we solve it for the case of zero $\dfrac{\partial\Delta k}{\partial h_{00}}=0$, we find the minority critical size:
$f_{0}^{*}=\dfrac{2+N}{4N}.$
This critical fraction depends on the network size $N$, but as the network size grows large, it approaches a fixed value:
$\lim_{N\to\infty}\dfrac{2+N}{4N}=\dfrac{1}{4}.$
Thus, in large networks, the critical minority size is approximately 25%.
Acknowledgments: We thank Jan Korbel for helpful feedback on our results. Competing interests: The author declare no competing interests.